# PDLP: A Practical First-Order Method for Large-Scale Linear Programming

### v1

@@ Abstract
To evaluate our method, we introduce a new collection of eleven large-scale LP problems with sizes ranging from 125 million to 6.3 billion nonzeros. PDLP solves eight of these instances to optimality gaps of 1% (with primal and dual feasibility errors of less than $10^{-8}$) within six days on a single machine. We also compare PDLP with Gurobi barrier, primal simplex, and dual simplex implementations. Gurobi barrier solves only three instances, exceeding our 1TB RAM limit on the other eight. While primal and dual simplex are more memory-efficient than the barrier method, they are slower and solve only three instances within six days.
Compared with the conference version of this work (in: Advances in Neural Information Processing Systems 34 (NeurIPS 2021)), the key new contributions are: (i) feasibility polishing, a technique that quickly finds solutions that are approximately optimal but almost exactly feasible (without which only three of the eleven problems can be solved); (ii) a multithreaded C++ implementation available in Google OR-Tools; and (iii) a new collection of large-scale LP problems. Note that the conference version should be referred to for comparisons with SCS and ablation studies, which we do not repeat in this paper.
@@ 1 Introduction
Despite their appealing computational properties, a longstanding barrier to adopting first-order methods for LP is their inability to reliably solve instances to high numerical precision in reasonable timeframes. For LP, it’s especially critical to find feasible solutions, both primal and dual. Even small violations of primal constraints (for instance, flow conservation) are difficult for practitioners to interpret. Similarly, finding dual feasible solutions is just as essential because of LP’s use for providing bounds on the optimal objective value such as within branch-and-bound methods [85]. As a concrete example of the risks associated with an inaccurately solved LP problem, consider the ‘irish-electricity’ instance, an LP instance based on a unit commitment case study of the Irish electric grid [63, 23], which has been used in Mittelmann’s LP benchmark [64]. On this instance, SCS 3.27 [72], a state-of-the-art first-order method with its default termination tolerance of $10^{-4}$, terminates in 40 seconds. However, the objective value for the returned solution is $2.21\times 10^{6}$, which is off from the true optimal objective value [^1] Found using Gurobi with crossover enabled. of $2.56\times 10^{6}$ by more than 13%. On the other hand, SCS with a tighter termination tolerance of $10^{-8}$ does not terminate successfully within an hour. This issue—where moderate feasibility violations lead to such undesirable behavior, which is only detectable using solutions provided from a more accurate solver—justifies practitioners’ caution in adopting first-order methods for LP.
To tackle this reliability issue, we apply a series of enhancements to PDHG including: adaptive restarting, dynamic primal-dual step size selection, and diagonal preconditioning. We call this enhanced PDHG algorithm, PDLP (PDHG for LP). The impact of these enhancements was demonstrated in the conference version of this paper [7]. For example, for finding high accuracy solutions on a standard LP benchmark, NETLIB [37], our enhancements enabled a 75×75$\times$75 × speed-up over PDHG without any enhancements and 19×19$\times$19 × speed-up [7, Table 24] over SCS [72], a first-order method for conic optimization.
@@
    A new collection of large-scale LP instances with between 125 million and 6.3 billion nonzeros to benchmark our method and a set of experiments demonstrating favorable results compared to Gurobi, a state of the art commerical solver. In comparison, the conference paper focused on smaller standard benchmark instances: the largest problem tested had 90 million nonzeros.
@@
Note that this paper does not subsume the conference version because we do not repeat its numerical experiments. These experiements were:
@@
    An ablation study that individually disabled each enchancement and also compared with reasonable alternatives.
@@
    Starting from a naive PDHG implementation, a demonstration of the incremental impact of each enchancement.
@@
    Comparisions with existing solvers (SCS and Gurobi) on instances from medium-size LP benchmarks and a large-scale LP example derived from the PageRank problem [19].
Given the C++ code is the exclusive focus of this paper, we decided neither to repeat results verbatim from the conference paper nor rerun these experiments using the C++ code as the conclusions would largely be the same. For reference [56, Table 8] shows that the single threaded C++ code achieves roughly a 1.5×1.5$\times$1.5 × speed-up over the Julia code.
@@ 1.1 Related literature
The conference version of this work [7] introduced PDLP, a restarted PDHG with several practical enhancements specialized for solving LP problems. Experiments were based on a prorotype Julia implementation. This paper includes an additional enhancement, and the experiments discussed in this paper are based on a C++ implementation of PDLP, which has been open-sourced through Google OR-Tools since 2022 [74]. More recently, a GPU implementation, dubbed cuPDLP, demonstrated strong numerical performance comparable to commercial LP solvers, garnering attention from both the academic community and the solver industry [56, 60]. Variants of (cu)PDLP have been incorporated into both commercial and open-source optimization solvers, including COPT [38], FICO Xpress [88], HiGHS [49], and NVIDIA cuOpt [16].
The success of PDLP has spurred theoretical analysis. Lu and Yang [55] proved that PDHG exhibits a linear convergence rate when applied to LP. Applegate et al. [8] showed that a faster linear convergence rate can be achieved through restarts, with the rate matching the worst-case lower bound. Applegate et al. [9] demonstrated that for infeasible or unbounded LP instances, PDHG iterates diverge toward a certificate of infeasibility or unboundedness, allowing for infeasibility detection without additional computation. Lu and Yang [58] established that PDHG displays a two-stage behavior when applied to LP: in the first stage, it identifies the optimal active variables in only finitely many iterations; in the second stage, it solves a homogeneous linear inequality system at an improved linear rate. This two-stage behavior helps explain the slow convergence observed when an LP instance is near-degenerate. Hinder [47] shows an iteration complexity bound for restarted PDHG applied to totally unimodular LP of $\tilde{O}(Hm^{2.5}\sqrt{\text{nnz}(A)})$ where $H$ is the largest absolute value of the right-hand side and objective vector, $m$ is the number of rows and $\text{nnz}(A)$ the number of nonzeros in the constraint matrix. Unlike other results, this bound does not involve any condition numbers. Zikai and Freund [87] introduced the concept of level-set geometry, showing that PDHG converges more quickly for LP instances with regular, non-flat level sets around the optimal solution. Recently, additional numerical improvements have been proposed to further accelerate the algorithm. In particular, Zikai and Freund [87] introduced a central-path-based preconditioner for PDHG that significantly improves convergence. Further, Lu and Yang [59] incorporated the Halpern iteration [43] into PDLP, achieving accelerated theoretical guarantees and improved practical performance.
Other first-order methods for LP
@@
  • ABIP and ABIP+ [54] are new interior-point method (IPM) solvers for cone programming. ABIP uses a few steps of ADMM to approximately solve the sub-problems that arise when applying a path-following barrier algorithm to the homogeneous self-dual embedding of the problem. Although IPMs are technically second-order methods, the use of ADMM in the inner loop allows the solver to scale to larger problems than traditional IPMs can handle. ABIP+ [30] adapts many of the improvements from PDLP to ABIP to improve its performance.
@@
  • ProxQP [10] is a solver specifically designed for convex quadratic programming (QP), particularly instances arising in real-time robotics applications. It uses a primal-dual proximal method of multipliers and requires solving a linear system at each iteration through Cholesky factorization. ProxQP can also leverage warm-start strategies, which is especially beneficial for robotics applications.
@@ 2 Preliminaries
Let ${\mathbb{R}}$ denote the set of real numbers, ${\mathbb{R}}^{+}$ the set of nonnegative real numbers, and ${\mathbb{R}}^{-}$ the set of nonpositive real numbers. Let ${\mathbb{N}}$ denote the set of natural numbers (starting from one). We endow ${\mathbb{R}}^{n}$ with the standard dot product $\langle x,y\rangle=x^{\top}y$ and its induced norm $\|\cdot\|_{2}.$ We will also use $\|\cdot\|_{p}$ to denote the $\ell_{p}$ norm for a vector. The $\ell_{\infty}$-norm, $\|v\|_{\infty}$, computes the maximum absolute value of the vector $v.$ For a vector $v\in{\mathbb{R}}^{n}$, we use $v^{+}$ and $v^{-}$ for their positive and negative parts, i.e., $v_{i}^{+}=\max\{0,v_{i}\}$ and $v_{i}^{-}=-\min\{0,v_{i}\}$. The symbol $v_{1:m}$ denotes the vector with the first $m$ components of $v$. The symbol $\mathbf{1}$ denotes the vector of all ones. Given a matrix $A\in{\mathbb{R}}^{m\times n},$ the symbols $A_{i,\cdot}$ and $A_{\cdot,j}$ correspond to the $i$th column and $j$th row of $A$, respectively. We use $\|\cdot\|_{2}$ to denote the spectral norm for a matrix, i.e., its top singular value.
@@
Given two vectors ({}$\ell\in({\mathbb{R}}\cup\{-\infty\})^{n}$ and $u\in({\mathbb{R}}\cup\{\infty\}\})^{n}$ satisfying $\ell\leq u.$ The function $p(\cdot;\ell,u)$ is piecewise linear and convex. With a slight abuse of notation, we will also use $p$ to denote an analogous function taking $m$-dimensional inputs as implied by the context.
@@
$\mathop{\rm minimize}_{x\in{\mathbb{R}}^{n}}~{}~{}&~{}c^{\top}x&\\ \text{subject to:}~{}~{}&~{}\ell_{c}\leq Ax\leq u_{c}\\ &~{}\ell_{v}\leq x\leq u_{v},\end{aligned}\qquad\qquad\qquad\begin{aligned} \mathop{\rm maximize}_{y\in{\mathbb{R}}^{m},r\in{\mathbb{R}}^{n}}\quad&-p(-y;\ell_{c},u_{c})-p(-r;\ell_{v},u_{v})\\ \text{subject to:}\quad&c-A^{\top}y=r\\ &y\in\mathcal{Y}\\ &r\in\mathcal{R}.\end{aligned}$ (1)
where $A\in{\mathbb{R}}^{m\times n}$, $c\in{\mathbb{R}}^{n}$, $\ell_{c}\in({\mathbb{R}}\cup\{-\infty\})^{m}$, $u_{c}\in({\mathbb{R}}\cup\{\infty\})^{m}$, $\ell_{v}\in({\mathbb{R}}\cup\{-\infty\})^{n}$, $u\in({\mathbb{R}}\cup\{\infty\})^{n}$, and the sets $\mathcal{Y}\subseteq{\mathbb{R}}^{m}$ and $\mathcal{R}\subseteq{\mathbb{R}}^{n}$ are Cartesian products with their $i$th components given by
$&\mathcal{Y}_{i}:=\begin{cases}\{0\}&(\ell_{c})_{i}=-\infty,~{}(u_{c})_{i}=\infty,\\ {\mathbb{R}}^{-}&(\ell_{c})_{i}=-\infty,~{}(u_{c})_{i}\in{\mathbb{R}},\\ {\mathbb{R}}^{+}&(\ell_{c})_{i}\in{\mathbb{R}},~{}(u_{c})_{i}=\infty,\\ {\mathbb{R}}&\text{otherwise};\end{cases}\quad\text{ and }\qquad\end{aligned}\begin{aligned} &\mathcal{R}_{i}:=\begin{cases}\{0\}&(\ell_{v})_{i}=-\infty,~{}(u_{v})_{i}=\infty,\\ {\mathbb{R}}^{-}&(\ell_{v})_{i}=-\infty,~{}(u_{v})_{i}\in{\mathbb{R}},\\ {\mathbb{R}}^{+}&(\ell_{v})_{i}\in{\mathbb{R}},~{}(u_{v})_{i}=\infty,\\ {\mathbb{R}}&\text{otherwise};\end{cases}\end{aligned}$ (2)
which corresponds with the set of variables $(y,r)$ for which the dual objective is finite. We highlight that the formulation we consider (1) is slightly different from the formulation in the conference version of this paper [7], which only considers a lower bound $\ell_{c}\leq Ax$ and has additional equality $Gx=q$.
@@
$\min_{x\in\mathcal{X}}\max_{y\in\mathcal{Y}}\mathcal{L}(x,y):=c^{\top}x+y^{\top}Ax-p(y;-u_{c},-\ell_{c})$ (3)
@@
When specialized to (3), the PDHG algorithm [24] takes the form:
$\begin{split}x^{k+1}&=\mathop{\bf proj}_{\mathcal{X}}(x^{k}-\tau(c-A^{\top}y^{k}))\\ y^{k+1}&=y^{k}-\sigma A\left(2x^{k+1}-x^{k}\right)-\sigma\mathop{\bf proj}_{[-u_{c},-\ell_{c}]}\left(\sigma^{-1}y^{k}-A\left(2x^{k+1}-x^{k}\right)\right)\end{split}$ (4)
where $\tau,\sigma>0$ are primal and dual step sizes, respectively. A quick case-by-case argument shows $y^{k+1}\in\mathcal{Y}$. For instance, suppose the $i$-th coordinate bounds satisfy $(\ell_{c})_{i}=-\infty$ and $(u_{c})_{i}\in\mathbb{R}$, and let $\widehat{y}^{k}\;=\;y^{k}\;-\;\sigma\,A\left(2\,x^{k+1}\;-\;x^{k}\right).$ Then, we have
@@
We call $\omega\in(0,\infty)$ the primal weight, and $\eta\in(0,\infty)$ the step size. Under this reparameterization, PDHG converges for all $\eta<1/\|A\|_{2}$. This parameterization allows us to control the scaling between the primal and dual iterates with a single parameter $\omega$. From now on and for the rest of the paper, we will use $z$ as a placeholder for the concatenation of primal dual vectors $(x,y).$ We use the term primal weight to describe $\omega$ because it is the weight on the primal variables in the following norm:

### v2

@@ Abstract
To evaluate our method, we introduce a new collection of eleven large-scale LP problems with sizes ranging from 125 million to 6.3 billion nonzeros. PDLP solves eight of these instances to optimality gaps of 1% (with primal and dual feasibility errors of less than $10^{-8}$) within six days on a single machine. We also compare PDLP with Gurobi barrier, primal simplex, and dual simplex implementations. These traditional methods are designed to solve linear programs to much tighter optimality gaps but struggle to solve these instances. Gurobi barrier solves only three instances, exceeding our 1TB RAM limit on the other eight. While primal and dual simplex are more memory-efficient than the barrier method, they are slower and solve only three instances within six days.
Compared with the conference version of this work (in: Advances in Neural Information Processing Systems 34 (NeurIPS 2021)), the key new contributions are: (i) feasibility polishing, a technique that quickly finds solutions that are approximately optimal but almost exactly feasible (without which only two of the eleven problems can be solved); (ii) a multithreaded C++ implementation available in Google OR-Tools; and (iii) a new collection of large-scale LP problems. Note that the conference version should be referred to for comparisons with SCS and ablation studies, which we do not repeat in this paper.
@@ 1 Introduction
Despite their appealing computational properties, a longstanding barrier to adopting first-order methods for LP is their inability to reliably solve instances to high numerical precision in reasonable timeframes. For LP, it is especially critical to find feasible solutions, both primal and dual. Even small violations of primal constraints (for instance, flow conservation) are difficult for practitioners to interpret. Similarly, finding dual feasible solutions is just as essential because of LP’s use for providing bounds on the optimal objective value such as within branch-and-bound methods [92]. As a concrete example of the risks associated with an inaccurately solved LP problem, consider the ‘irish-electricity’ instance, an LP instance based on a unit commitment case study of the Irish electric grid [70, 25], which has been used in Mittelmann’s LP benchmark [71]. On this instance, SCS 3.27 [79], a state-of-the-art first-order method with its default termination tolerance of $10^{-4}$, terminates in 40 seconds. However, the objective value for the returned solution is $2.21\times 10^{6}$, which is off from the true optimal objective value [^1] Found using Gurobi with crossover enabled. of $2.56\times 10^{6}$ by more than 13%. On the other hand, SCS with a tighter termination tolerance of $10^{-8}$ does not terminate successfully within an hour. This issue—where moderate feasibility violations lead to such undesirable behavior, which is only detectable using solutions provided from a more accurate solver—justifies practitioners’ caution in adopting first-order methods for LP.
To tackle this reliability issue, we apply a series of enhancements to PDHG including: adaptive restarting, dynamic primal-dual step size selection, and diagonal preconditioning. We call this enhanced PDHG algorithm, PDLP (PDHG for LP). The impact of these enhancements was demonstrated in the conference version of this paper [9]. For example, for finding high accuracy solutions on a standard LP benchmark, NETLIB [42], our enhancements enabled a $75\times$ speed-up over PDHG without any enhancements and a $19\times$ speed-up [9, Table 24] over SCS [79], a first-order method for conic optimization.
@@
    A new collection of large-scale LP instances with between 125 million and 6.3 billion nonzeros to benchmark our method and a set of experiments demonstrating favorable results compared to Gurobi, a state of the art commercial solver. In comparison, the conference paper focused on smaller standard benchmark instances: the largest problem tested had 90 million nonzeros.
@@
Note that this paper does not subsume the conference version because we do not repeat its numerical experiments. These experiments were:
@@
    An ablation study that individually disabled each enhancement and also compared with reasonable alternatives.
@@
    Starting from a naive PDHG implementation, a demonstration of the incremental impact of each enhancement.
@@
    Comparisons with existing solvers (SCS and Gurobi) on instances from medium-size LP benchmarks and a large-scale LP example derived from the PageRank problem [21].
Given the C++ code is the exclusive focus of this paper, we decided neither to repeat results verbatim from the conference paper nor rerun these experiments using the C++ code as the conclusions would largely be the same. For reference [66, Table 8] shows that the single threaded C++ code achieves roughly a $1.5\times$ speed-up over the Julia code.
@@ 1.1 Related literature
There have also been several papers that modify interior point methods so that instead of using a factorization to solve the linear systems that arise at each iteration they use preconditioned iterative methods [47, 40, 98, 48, 33]. This reduces the memory overhead of these methods and the cost of each inner iteration performed by the iterative method (e.g., conjugate gradient) is roughly the cost of a matrix-vector multiplication (assuming the preconditioner is cheap to implement). For example, Fountoulakis et al. [40] demonstrate excellent performance of these methods on large-scale compressed sensing problems.
@@
The conference version of this work [9] introduced PDLP, a restarted PDHG with several practical enhancements specialized for solving LP problems. Experiments were based on a prototype Julia implementation. This paper includes an additional enhancement, and the experiments discussed in this paper are based on a C++ implementation of PDLP, which has been open-sourced through Google OR-Tools since 2022 [81]. More recently, a GPU implementation, dubbed cuPDLP, demonstrated strong numerical performance comparable to commercial LP solvers, which garnered attention from both the academic community and the solver industry [66, 67]. Variants of (cu)PDLP have been incorporated into both commercial and open-source optimization solvers, including COPT [43], FICO Xpress [95], HiGHS [56], and NVIDIA cuOpt [18].
The success of PDLP has spurred theoretical analysis. Lu and Yang [62] proved that PDHG exhibits a linear convergence rate when applied to LP. Applegate et al. [10] showed that a faster linear convergence rate can be achieved through restarts, with the rate matching the worst-case lower bound. Applegate et al. [11] demonstrated that for infeasible or unbounded LP instances, PDHG iterates diverge toward a certificate of infeasibility or unboundedness, allowing for infeasibility detection without additional computation. Lu and Yang [65] established that PDHG displays a two-stage behavior when applied to LP: in the first stage, it identifies the optimal active variables in finitely many iterations; in the second stage, it solves a homogeneous linear inequality system at an improved linear rate. This two-stage behavior helps explain the slow convergence observed when an LP instance is near-degenerate. Hinder [54] shows an iteration complexity bound for restarted PDHG applied to totally unimodular LP of $O(Hm^{2.5}\sqrt{\text{nnz}(A)}\ln(Hn/\epsilon))$ where $H$ is the largest absolute value of the right-hand side and objective vector, $n$ is the number of columns, $m$ is the number of rows, $\text{nnz}(A)$ the number of nonzeros in the constraint matrix, and $\epsilon$ is the distance to optimality measured in Euclidean norm. Unlike other results, this bound does not involve any condition numbers. Xiong and Freund [94] introduced the concept of level-set geometry, showing that PDHG converges more quickly for LP instances with regular, non-flat level sets around the optimal solution. Recently, additional numerical improvements have been proposed to further accelerate the algorithm. In particular, Xiong and Freund [94] introduced a central-path-based preconditioner for PDHG that significantly improves convergence. Further, Lu and Yang [64] incorporated the Halpern iteration [50] into PDLP, achieving accelerated theoretical guarantees and improved practical performance.
Other first-order methods for LP.
@@
  • ABIP and ABIP+ [61] are interior-point method (IPM) solvers for cone programming. ABIP uses a few steps of ADMM to approximately solve the sub-problems that arise when applying a path-following barrier algorithm to the homogeneous self-dual embedding of the problem. Although IPMs are technically second-order methods, the use of ADMM in the inner loop allows the solver to scale to larger problems than traditional IPMs can handle. ABIP+ [34] adapts many of the improvements from PDLP to ABIP to improve its performance.
@@
  • ALGENCAN [7] is a widely established software package for nonlinear programming that is based on the augmented Lagrangian method. The augmented Lagrangian subproblems are solved by the inner solver GENCAN [17], which utilizes active-set strategies and spectral projected gradients. Crucially, ALGENCAN is designed to operate in a fully matrix-free manner, i.e., only using gradient evaluations and Hessian-vector products.
@@ 2 Preliminaries
Let $\mathbb{R}$} denote the set of real numbers, $\mathbb{R}$$^{+}$ the set of nonnegative real numbers, and $\mathbb{R}$$^{-}$ the set of nonpositive real numbers. Let $\mathbb{N}$} denote the set of natural numbers (starting from one). We endow ${\mathbb{R}}^{n}$ with the standard dot product $\langle x,y\rangle=x^{\top}y$ and its induced norm $\|\cdot\|_{2}.$ We will also use $\|\cdot\|_{p}$ to denote the $\ell_{p}$ norm for a vector. The $\ell_{\infty}$-norm, $\|v\|_{\infty}$, computes the maximum absolute value of the vector $v.$ For a vector $v\in{\mathbb{R}}^{n}$, we use $v^{+}$ and $v^{-}$ for their positive and negative parts, i.e., $v_{i}^{+}=\max\{0,v_{i}\}$ and $v_{i}^{-}=-\min\{0,v_{i}\}$. The symbol $\mathbf{1}$ denotes the vector of all ones. Given a matrix $A\in{\mathbb{R}}^{m\times n},$ the symbols $A_{i,\cdot}$ and $A_{\cdot,j}$ correspond to the $i$th row and $j$th column of $A$, respectively. We use $\|\cdot\|_{2}$ to denote the spectral norm for a matrix, i.e., its top singular value. We use $z^{k}=(x^{k},y^{k})$ to denote the $k$th iterate of PDHG and $z^{n,t}=(x^{n,t},y^{n,t})$ to denote iterates of PDLP where $n$ is the outer iteration counter and $t$ is the inner iteration counter.
@@
The function $p(\cdot;\ell,u)$ is piecewise linear and convex for any two vectors $\ell\in({\mathbb{R}}\cup\{-\infty\})^{n}$ and $u\in({\mathbb{R}}\cup\{\infty\})^{n}$ satisfying $\ell\leq u$. With a slight abuse of notation, we will also use $p$ to denote an analogous function taking $m$-dimensional inputs as implied by the context.
@@
$\begin{aligned} \mathop{\rm minimize}_{x\in{\mathbb{R}}^{n}}~~&~c^{\top}x&\\ \text{subject to:}~~&~\ell_{c}\leq Ax\leq u_{c}\\ &~\ell_{v}\leq x\leq u_{v},\end{aligned}\qquad\qquad\qquad\begin{aligned} \mathop{\rm maximize}_{y\in{\mathbb{R}}^{m},r\in{\mathbb{R}}^{n}}\quad&-p(-y;\ell_{c},u_{c})-p(-r;\ell_{v},u_{v})\\ \text{subject to:}\quad&c-A^{\top}y=r\\ &y\in\mathcal{Y}\\ &r\in\mathcal{R}.\end{aligned}$ (1)
where $A\in{\mathbb{R}}^{m\times n}$, $c\in{\mathbb{R}}^{n}$, $\ell_{c}\in({\mathbb{R}}\cup\{-\infty\})^{m}$, $u_{c}\in({\mathbb{R}}\cup\{\infty\})^{m}$, $\ell_{v}\in({\mathbb{R}}\cup\{-\infty\})^{n}$, $u_{v}\in({\mathbb{R}}\cup\{\infty\})^{n}$, and the sets $\mathcal{Y}\subseteq{\mathbb{R}}^{m}$ and $\mathcal{R}\subseteq{\mathbb{R}}^{n}$ are Cartesian products with their $i$th components given by
$\begin{aligned} &\mathcal{Y}_{i}:=\begin{cases}\{0\}&(\ell_{c})_{i}=-\infty,~(u_{c})_{i}=\infty,\\ {\mathbb{R}}^{-}&(\ell_{c})_{i}=-\infty,~(u_{c})_{i}\in{\mathbb{R}},\\ {\mathbb{R}}^{+}&(\ell_{c})_{i}\in{\mathbb{R}},~(u_{c})_{i}=\infty,\\ {\mathbb{R}}&\text{otherwise};\end{cases}\quad\text{ and }\qquad\end{aligned}\begin{aligned} &\mathcal{R}_{i}:=\begin{cases}\{0\}&(\ell_{v})_{i}=-\infty,~(u_{v})_{i}=\infty,\\ {\mathbb{R}}^{-}&(\ell_{v})_{i}=-\infty,~(u_{v})_{i}\in{\mathbb{R}},\\ {\mathbb{R}}^{+}&(\ell_{v})_{i}\in{\mathbb{R}},~(u_{v})_{i}=\infty,\\ {\mathbb{R}}&\text{otherwise};\end{cases}\end{aligned}$
which are the values of $y$ and $r$ with a finite dual objective value. We highlight that the formulation we consider (1) is slightly different from the formulation in the conference version of this paper [9], which only considers a lower bound $\ell_{c}\leq Ax$ and has additional equality constraints $Gx=q$.
@@
$\min_{x\in\mathcal{X}}\max_{y\in\mathcal{Y}}\mathcal{L}(x,y):=c^{\top}x-y^{\top}Ax-p(y;-u_{c},-\ell_{c})$ (2)
@@
When specialized to (2), the PDHG algorithm takes the form:
$x^{k+1}$ $=\mathop{\bf proj}_{\mathcal{X}}(x^{k}-\tau(c-A^{\top}y^{k}))$ (3a)
$y^{k+1}$ $=y^{k}-\sigma A\left(2x^{k+1}-x^{k}\right)-\sigma\mathop{\bf proj}_{[-u_{c},-\ell_{c}]}\left(\sigma^{-1}y^{k}-A\left(2x^{k+1}-x^{k}\right)\right)$ (3b)
where $\tau,\sigma>0$ are primal and dual step sizes, respectively. For completeness we provide the derivation of Equation 3 from PDHG in Appendix B. A quick case-by-case argument shows $y^{k+1}\in\mathcal{Y}$. For instance, suppose the $i$-th coordinate bounds satisfy $(\ell_{c})_{i}=$-$\infty$ and $(u_{c})_{i}\in\mathbb{R}$, and let $\widehat{y}^{k}\;=\;y^{k}\;-\;\sigma\,A\left(2\,x^{k+1}\;-\;x^{k}\right).$ Then, we have
@@
We call $\omega\in(0,\infty)$ the primal weight, and $\eta\in(0,\infty)$ the step size. Under this reparameterization, PDHG converges for all $\eta<1/\|A\|_{2}$. This parameterization allows us to control the scaling between the primal and dual iterates with a single parameter $\omega$. From now on and for the rest of the paper, we will use $z$ as a placeholder for the concatenation of primal and dual vectors $(x,y).$ We use the term primal weight to describe $\omega$ because it is the weight on the primal variables in the following norm: